OpenStudy (anonymous):

Name Melting Point Density (g/cm3) Zinc 419.53 oC 7.134 Nickel 1455oC 8.912 Copper 1084.62oC 8.96 Silver 961.78oC 10.501 Gold 1064.18oC 19.282 Determine the identity of a cube of metal that measures 1.5 cm on each side and has a mass of 35.45g. A) Gold B) Nickel C) Silver D) Zinc

OpenStudy (shikamaru11):

Silver :) Bet it divide the mass of each metal by given volume so the close one is the answer @beyonced98

OpenStudy (shikamaru11):

that's for silver 107/3 = 35.66g

OpenStudy (michele_laino):

volume of the cube is: \(V=1.5^3=...cm^3\) so density is: \(\delta=m/V= 35.45/ V=...?\)

OpenStudy (shikamaru11):

@Michele_Laino Sir what mass we should put to achieve the answer i.e given 35.45 i think that's what he asks

OpenStudy (michele_laino):

you have to use \(m=35.45\; grams\)

OpenStudy (shikamaru11):

but Sir @Michele_Laino it's given

OpenStudy (shikamaru11):

Got it Sir THAnKS